OpenStudy (anonymous):

find the consecutive integers m and n such that given number is between m and n or the given number is not a real number. -√-90

OpenStudy (radar):

I take it that the given number is \[-\sqrt{-90}\]

OpenStudy (anonymous):

yes @radar

OpenStudy (radar):

Well the square root of -90 is an imaginary number it could be expressed ans\[-\sqrt{-9(10)}=-3i \sqrt{10}\] still an imaginary number, therefore not part of the real number definition.

OpenStudy (anonymous):

so you mean all number that is negative in square root is an imaginary number?

OpenStudy (radar):

No, I mean the roots of square roots of negative numbers is imaginary, the result of the radical will lead to imaginary (not real) numbers.

OpenStudy (anonymous):

@radar thanks a lot!

OpenStudy (radar):

It is kind quirky, the letter i is used to indicate the square root of -1 so for example:\[\sqrt{-4}=2i \] an imaginary number. In electronics the i is referred as the "j" operator.

OpenStudy (anonymous):

Thanks a lot @radar

OpenStudy (radar):

You are most welcome.